OpenStudy (anonymous):

Please help with this problem. I spent 2 hours working in this problem. Nobody in my class got it, so if you get it, you are the smartest person I know. if f(x) = 3/(6pi - 9x) Then the slope of the tangent line to the graph of y=f(x) at the point (pi,-3/3pi) is the limit as x tends to ______ of the following expression ___________. The Value of the limit is _________ It follows that the equation of the tangent line is y = _____________. Please help if you can thank you.

OpenStudy (dumbcow):

slope of tangent line is the derivative of the function at the given point ---> f'(pi) use chain rule, Let u = 6pi - 9x , du = -9 \[f'(x) = \frac{3}{u} *du = -\frac{3}{u^2}*du = \frac{27}{(6\pi -9x)^2}\] pluggin in x=pi \[f'(\pi) = \frac{27}{(-3\pi)^2} = \frac{3}{\pi^2}\]

OpenStudy (anonymous):

\[f'(x) = \lim_{x \rightarrow a}\frac{ f(x)-f(a) }{ x-a }\] ... you want to find the value of this limit (as x tends to the x-coordinate of the given point (a, f(a)) and the expression is what you get when you fill in your given function for f(x) and your given x-coordinate for a and given y-coordinate for f(a). to then find the equation of the tangent line, you want to use the value of the limit as x approaches a (that gives you the slope of the tangent line) and the point given with y=mx+b to get the value of b... then you've go tthe equation of the tangent line